OpenStudy (anonymous):

If sin Θ =5/6 , what are the values of cos Θ and tan Θ

OpenStudy (anonymous):

are you supposed to use inverse sine for this?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

do you know how to solve it

OpenStudy (anonymous):

yes, we know the identity sin^2 x + cos^x = 1 so solve that equation for cos^2 x and we get cos^2 x = 1 - sin^2x then taking the square root of both sides cos x = sqrt(1 - sin^2 x) we know sin x = 5/6 so substitute that in and simplify